# Invisible Walls in Cities: Leveraging Large Language Models to Predict Urban Segregation Experience with Social Media Content

### v2

@@ 7.1 Experiment Settings
We employ GPT-4o for reflective coding on a small subset (n=190) of POIs, and GPT-4o mini for codebook-guided reasoning on all POIs for economic considerations. All models are trained using the Adam optimizer to minimize the MSE loss, with hyperparameters (e.g., learning rate, weight decay) tuned based on validation performances. POIs in each city are randomly split into training, validation, and test sets with a ratio of 6:2:2. For further implementation details, please refer to Appendix C and our released code [^5] https://anonymous.4open.science/r/REEM-73DB/.
Model MSE↓ RMSE↓ MAE↓ R2↑
@@
City POI Num Model MSE↓ RMSE↓ MAE↓ R2↑
@@ 7.2 Experiment Results
To validate our design of different model components, we perform an ablation study. As Table 3 shows, each adapter (Reasoning, Embedding, Population) exhibits certain predictive capabilities, but none achieves optimal performance. The full RE’EM model, which strategically combines all three adapters, consistently delivers superior results. Thus, the integration of multi-view signals enables complementary information flow, allowing the model to refine its predictions beyond what any single feature source can achieve.
@@
8 Implication
Our work highlights the transformative potential of LLMs to uncover segregation experiences encoded in social media content. By revealing the nuanced and often invisible barriers that shape social interactions from online reviews, this research exemplifies the ethical and socially beneficial application of AI, demonstrating how advanced technologies can be harnessed to foster inclusivity and promote more equitable urban environments. Policymakers, urban planners, and community leaders can leverage these insights to better understand and address patterns of social exclusion, ultimately building more cohesive and diverse communities.
From a technical perspective, our framework contributes to the application of LLMs for tackling complex societal challenges. With reflective coding and integration of LLMs’ reasoning and embedding capabilities, our framework moves beyond conventional predictive tasks to offer actionable insights. Furthermore, our work demonstrates how prompt-guided analytical processes can mitigate biases and hallucinations, paving the way for more reliable and socially responsible AI applications.
@@ 9 Conclusion
In this paper, we pioneer the use of social media data to predict experienced segregation, designing an LLM Reflective Coder to generate insightful summaries and a REasoning-and-EMbedding (RE’EM) framework to integrate reasoning and embedding for accurate predictions. Our approach is validated through both user studies and quantitative experiments, demonstrating its effectiveness in producing human-comprehensible review summaries and reliable segregation predictions. This work not only advances the understanding of experienced segregation but also provides a foundation for leveraging social media data to address broader societal challenges, such as fostering inclusivity and mitigating social inequalities.
Ethical Statement
The Yelp review data was collected under users’ consent and fully anonymized. The Safegraph mobility data was aggregated to the CBG level by month and processed with differential privacy techniques to safeguard against individual identification. The ACS demographic data is publicly accessible at the CBG level. Thus, no approval from the Institutional Review Board was required by the authors’ institutions.
@@ Appendix B Reflective Coding Algorithm
1:Input: A set $O$ of tuple (POI $I$, POI name $N$, POI reviews and images $M$, real visitor composition $\tau_{real}$, local residential composition $l_{res}$)
@@
4:for each tuple ($I$, $N$, $M$, $tau_{real}$, $l_{res}$) in $O$ do
@@ Appendix D Case study
Figure 6: Case study.

### v3

@@ 7.1 Experiment Settings
We employ GPT-4o for reflective coding on a small subset (n=190) of POIs, and GPT-4o mini for codebook-guided reasoning on all POIs for economic considerations. All models are trained using the Adam optimizer to minimize the MSE loss, with hyperparameters (e.g., learning rate, weight decay) tuned based on validation performances. POIs in each city are randomly split into training, validation, and test sets with a ratio of 6:2:2. For further implementation details, please refer to Appendix C and our released code  [^5] https://github.com/LinChen-65/REEM.
Model MSE↓ RMSE↓ MAE↓ R2
@@
City POI Num Model MSE↓ RMSE↓ MAE↓ R2
@@ 7.2 Experiment Results
To validate our design of different model components, we perform an ablation study. As Table 3 shows, each channel (reasoning, embedding, population) exhibits certain predictive capabilities, but none achieves optimal performance. The full RE’EM model, which strategically combines all three channels, consistently delivers superior results. Thus, the integration of multi-view signals enables complementary information flow, allowing the model to refine its predictions beyond what any single feature source can achieve.
@@
8 Discussion
Our work highlights the transformative potential of LLMs in uncovering segregation experiences encoded in social media content. By revealing the nuanced and often invisible barriers that shape social interactions from online reviews, our work exemplifies the ethical and socially beneficial application of AI, demonstrating how advanced technologies can be harnessed to foster inclusivity and promote more equitable urban environments. Policymakers, urban planners, and community leaders can leverage these insights to better understand and address patterns of social exclusion, ultimately working toward the development of more cohesive and diverse communities.
From a technical perspective, our framework advances the application of LLMs in tackling complex societal challenges. With reflective coding and integration of LLMs’ reasoning and embedding capabilities, our framework not only significantly improves prediction accuracy but extends beyond conventional predictive tasks to offer explanations and actionable insights. Furthermore, our work demonstrates how prompt-guided analytical processes can mitigate biases and hallucinations, paving the way for more reliable and socially responsible AI applications.
Our work has several limitations. Despite rigorous efforts to filter POIs with adequate review coverage, inherent imbalances in demographic representation across social media platforms may skew corpus distributions. While our work demonstrates the overall effectiveness of using social media data to predict segregation, future work could further explore how and to what extent these data biases impact prediction accuracy across different POIs. Besides, our analysis does not account for temporal variations in segregation, such as those driven by POI updates, policy shifts, or acute events like pandemics and economic crises. Future work could integrate temporal modeling to capture these dynamics.
@@ 9 Conclusion
In this paper, we pioneer the use of social media data to predict experienced segregation, designing a reflective LLM coder to generate insightful summaries and a REasoning-and-EMbedding (RE’EM) framework to integrate reasoning and embedding for accurate predictions. Our approach is validated through a qualitative user study and quantitative experiments, demonstrating its effectiveness in producing human-comprehensible review summaries and reliable segregation predictions. This work not only advances the understanding of experienced segregation but also provides a foundation for leveraging social media data to address broader societal challenges, such as fostering inclusivity and mitigating social inequalities.
@@ Appendix B Reflective Coding Algorithm
1:Input: A set $O$ of tuple (POI II, POI name NN, POI reviews and images $M$, real visitor composition $\tau_{real}$, local residential composition $l_{res}$)
@@
4:for each tuple (II, NN, $M$, $tau_{real}$, $l_{res}$) in $O$ do
@@ Appendix C Implementation Details
During the development of REEM, we conduct a compact grid search over the principal hyper-parameters using the Philadelphia dataset. Specifically, learning rates for both the GAT and fully-connected components are explored in the interval $5\times 10^{-6},\ 5\times 10^{-3}]$, weight-decay coefficients are varied from $1\times 10^{-5}$ to $5\times 10^{-4}$, and the early-stopping patience is examined between $10$ and $25$ epochs. The width of every MLP hidden layer is tuned by doubling or halving around the final choices (approximately 6464512512 units). The configuration achieving the lowest validation MSE is selected and is released as the default value in our open-source code.
All experiments are conducted on a Linux server with the following hardware and software specifications:
  • CPU: Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz
[…]
  • Memory: 503.5 GB
  • Operating System: Ubuntu 22.04.3 LTS
  • Software:
    • Python 3.11.5
    • PyTorch 2.4.0 (with CUDA 12.1 support)
    • Other libraries: NumPy 1.26.0, vllm 0.6.1.post2
@@ Appendix D Case study
Appendix E Ethics Statement
All datasets in this study were sourced from publicly available or academically licensed repositories, with privacy protections in place. The review data was obtained from the Yelp Open Dataset under its Data Licensing agreement. The data was collected under users’ consent and fully anonymized. The Safegraph mobility data is currently accessible through the Dewey Data platform [^6] https://www.deweydata.io/ for academic purposes. This data was aggregated to the CBG level by month and processed with differential privacy techniques to safeguard against individual identity leakage. The ACS demographic data is publicly available [^7] https://www.census.gov/programs-surveys/acs/. As such, no Institutional Review Board (IRB) approval was required by the authors’ institutions.
Appendix F Code and Data Availability
All source code used in this work, including LLM prompts, hyperparameter configurations, data preprocessing scripts, and training scripts, is available [^8] https://github.com/LinChen-65/REEM. Given that the datasets are publicly accessible and of substantial size, we provide detailed instructions for accessing the required data, rather than including the raw data directly.